Find the electrostatic force between charges of +2.0 C and +5.0 C separated by a distance of 75m in a vacuum.

1
Expert's answer
2022-02-13T15:14:43-0500

Given:

"q_1=\rm 2.0\:C"

"q_2=\rm 5.0\:C"

"r=\rm 75\:\rm m"


The Coulomd law says


"F=k\frac{q_1q_2}{r^2}"

Hence


"F=9*10^9*\frac{2.0*5.0}{75^2}=1.6*10^7\:\rm N"